OpenStudy (anonymous):

calculate the number whose difference between itself and its square is a maximum

OpenStudy (anonymous):

Incomplete info. as it is x-x^2=max

OpenStudy (anonymous):

how is the info incomplete?

OpenStudy (anonymous):

a maximum: a maximum means nothing in of itself, a maximum to what?!

OpenStudy (anonymous):

the maximum difference between any two real numbers

OpenStudy (anonymous):

sorry the maximmum difference between any number and its square

OpenStudy (anonymous):

for example for 0.8 the difference is .8 - .8^2 = .16

OpenStudy (watchmath):

Ok, so we want to maximize \(f(x)=x-x^2\) Without calculus: find the vertex \(x=-b/2a=-1/(2(-1))=1/2\). So it is maximum when \(x=1/2). With calculus, set \(f'(x)=1-2x=0\). Hence \(x=1/2\). Then you just need to justify that x=1/2 gives the maximum value.

OpenStudy (anonymous):

this is how i interpreted the question

OpenStudy (anonymous):

right! how the heck didn't i think of that - ty. how do i give medals?

OpenStudy (watchmath):

Press the good answer button :)